1. Show whether or not the following differential equations are separable:

1.1 "\frac{dy}{dx}=\frac{x+1}{y-1}"


1.2 "\frac{dy}{dx}=\frac{ye^{x}+y}{x^2+2}"



1.3 "\frac{ds}{dt}=t\ln(s^{2t})+8t^2"


1
Expert's answer
2021-08-13T08:15:14-0400

1.1


"\dfrac{dy}{dx}=\dfrac{x+1}{y-1}"

"(y-1)dy=(x+1)dx"

The equation is separable.


1.2


"\dfrac{dy}{dx}=\dfrac{ye^x+y}{x^2+2}"

"\dfrac{dy}{dx}=y(\dfrac{e^x+1}{x^2+2})"

"\dfrac{1}{y}dy=(\dfrac{e^x+1}{x^2+2})dx"

The equation is separable.


1.3


"\dfrac{ds}{dt}=t\ln(s^{2t})+8t^2"

"\dfrac{ds}{dt}=t(2t)\ln(s)+8t^2"

"\dfrac{ds}{dt}=2t^2(\ln(s)+4)"




"\dfrac{ds}{\ln(s)+4}=2t^2dt"

The equation is separable.